# Semi-Riemannian $\text{spin}^{c}$ manifolds carrying generalized Killing spinors and the classification of Riemannian $\text{spin}^{c}$ manifolds admitting a type I imaginary generalized Killing spinor

### v1

@@ 1 Introduction
The above results have been extended by many in various directions. Among these, the following classifications are directly related to this article. Moroianu [34] classifies complete, simply connected Riemannian $\text{spin}^{c}$ manifolds carrying parallel spinors. Moroianu shows that the manifold splits as a Riemannian product between a Kähler manifold (equipped with its canonical $\text{spin}^{c}$-structure) and a spin manifold carrying a parallel spinor [34, Theorem 1.1]. Shortly after Baum, Rademacher extends Baum’s results to what we call special imaginary generalized Killing spinors [36]. Große and Nakad have extended Rademacher’s classification to the $\text{spin}^{c}$ setting [22, Theorem 4.1]. An extension of Rademacher’s result to the setting of Riemannian spin manifolds carrying a type I imaginary generalized Killing spinor have been made by Leistner and Lischewski [32, Theorem 4].
@@ 2 Preliminaries and notation
This section is a summary of some preliminaries from spin and $\text{spin}^{c}$ geometry. The reader is directed to [30] and [16] for details.
@@
Let $(M,g)$ be a time oriented semi-Riemannian $\text{spin}^{c}$ manifold of dimension $n$ and index $r$. From an irreducible complex representation $\sigma$ of the $\text{spin}^{c}$ group, we obtain the complex spinor bundle $S\to M$, which is a vector bundle of complex dimension $2^{\lfloor\frac{n}{2}\rfloor}$. When $n$ is even, $S$ splits as a direct sum $S=S^{+}\oplus S^{-}$.
@@
where $\{e_{1},\ldots,e_{n}\}$ is a pseudo-orthonormal frame with $\varepsilon_{k}=g(e_{k},e_{k})$ and $i\omega\in\Omega^{2}(M,i\mathbb{R})$ denotes the curvature 2-form coming from the connection 1-form $A\in\Omega^{1}(P_{S^{1}},i\mathbb{R})$ [16, Page 65].
@@ 4.1 Non-local splittings without completeness assumptions
Fix some uniform pair $(I,L_{p}^{\prime})$ for $\Phi$. First, assume that $\Phi|_{I\times L_{p}^{\prime}}$ is injective. Take some point $q\in L_{p}^{\prime}$, with $q\neq p$. Since $L_{p}^{\prime}$ is connected, there exists a path with embedded image $\ell\cong(0,1)$ in $L_{p}^{\prime}$, containing both the point $p$ and the point $q$. Take a tubular neighborhood $\mathcal{U}$ of $\ell$ in $L_{p}^{\prime}$. Since $V$ is nowhere tangent to $\mathcal{U}$ and $\Phi|_{I\times\hskip 0.85358pt\mathcal{U}}$ is injective, we have that $\Phi|_{I\times\hskip 0.85358pt\mathcal{U}}$ is a diffeomorphism onto an open submanifold $N$ of $M$ [31, Theorem 9.20]. Since $N$ has trivial de Rahm cohomology, we can write $V|_{N}=\operatorname{grad}h$, for some function $h:N\to\mathbb{R}$ and note that the intersection of each leaf $L_{m}\in\mathcal{F}$ with $N$ is either empty or a level set of $h$. Hence, if we show that $h(\Phi_{t}(p))=h(\Phi_{t}(q))$, for all $t\in I$, then $\Phi_{t}(q)\in L_{\Phi_{t}(p)}$ and we could conclude the current case. By assumption, it is true for $t=0$. Therefore, it suffices to show that $\frac{d}{dt}h(\Phi_{t}(p))=\frac{d}{dt}h(\Phi_{t}(q))$ for all $t\in I$. We compute that
@@ 5.1 The semi-Riemannian $\text{spin}^{c}$ setting
In the Lorentzian case, we have stronger results due to Bohle [11, Section 5] and Leitner [33, Theorem 5.3]. In the Riemannian case, the global statement is essentially the same as due to Baum in [7]. The author is not aware of similar local statements in the Riemannian case but is sure that they are known to experts. There are also stronger results for $r=2$, due to Shafiee and Bahrampour [37]. In case of parallel spinors we have stronger results in [15, 10, 38, 26].
@@ 5.2.1 Type I imaginary generalized Killing spinors
In what follows, we will apply the $\text{spin}^{c}$ Gauß formula from [35, Section 3]. We follow the setup of that paper. Given a leaf $L_{p}\in\mathcal{F}$, we equip it with the induced $\text{spin}^{c}$ structure and we write $S_{L_{p}}$ to denote the induced spinor bundle on the hypersurface $L_{p}$. When the dimension $n$ of $M$ is odd, we have that $S_{L_{p}}\cong S|_{L_{p}}$ and when $n$ is even, we have that $S_{L_{p}}\cong S^{+}|_{L_{p}}$. Hence, when $n$ is odd, we can simply write $\psi|_{L_{p}}$ to denote the restriction of a section, but when $n$ is even we write $\psi|_{L_{p}}$, to mean $\psi^{+}|_{L_{p}}$. In this setup, the $S^{1}$-bundle associated with the $\text{spin}^{c}$ structure on $L_{p}$ is simply the restriction $P_{S^{1}}|_{L_{p}}$, where $P_{S^{1}}$ is the $S^{1}$-bundle associated to the $\text{spin}^{c}$ structure on $M$. Hence, given a connection 1-form $A\in\Omega^{1}(P_{S^{1}},i\mathbb{R})$, we get induced connections $\nabla^{S}$ on $S$ and $\nabla^{S_{L_{p}}}$ on $S_{L_{p}}$. If we let $\bullet:T(L_{p})\otimes S_{L_{p}}\to S_{L_{p}}$ denote the Clifford multiplication on $S_{L_{p}}$, we have that $X\bullet\eta=\tilde{E}\cdot X\cdot\eta$, where on the right hand side, $\eta$ is either viewed as an element of $S|_{L_{p}}$ or $S^{+}|_{L_{p}}$, depending on if $n$ is odd or even.
@@ 5.2.2 Type II imaginary generalized Killing spinors
Recall that an imaginary generalized Killing spinor on a Riemannian $\text{spin}^{c}$ manifold is said to be of type II if the constant $Q_{-}=\langle{\varphi,\varphi}\rangle^{2}-g(V_{\varphi},V_{\varphi})$ is positive. In case the type II imaginary generalized Killing spinor is a Killing spinor, then Große and Nakad [22, Proposition 4.3] showed that the manifold is isometric to hyperbolic space, provided it is complete. It was shown by Rademacher [36] that on spin manifolds, all type II special imaginary generalized Killing spinors have constant Killing function, i.e. they are Killing spinors. Grosse and Nakad [22, Section 4.2.2] extends this to the $\text{spin}^{c}$ case for dimension at least $3$ and provide a counterexample in dimension $2$.

### v2

@@ 1 Introduction
The above results have been extended by many in various directions. Among these, the following classifications are directly related to this article. Moroianu [35] classifies complete, simply connected Riemannian $\text{spin}^{c}$ manifolds carrying parallel spinors. Moroianu shows that the manifold splits as a Riemannian product between a Kähler manifold (equipped with its canonical $\text{spin}^{c}$-structure) and a spin manifold carrying a parallel spinor [35, Theorem 1.1]. Shortly after Baum, Rademacher extends Baum’s results to what we call special imaginary generalized Killing spinors [37]. Große and Nakad have extended Rademacher’s classification to the $\text{spin}^{c}$ setting [23, Theorem 4.1]. An extension of Rademacher’s result to the setting of Riemannian spin manifolds carrying a type I imaginary generalized Killing spinor have been made by Leistner and Lischewski [33, Theorem 4].
@@ 2 Preliminaries and notation
This section is a summary of some preliminaries from spin and $\text{spin}^{c}$ geometry. The reader is directed to [31, 14, 17] for details.
@@
Let $(M,g)$ be a time oriented semi-Riemannian $\text{spin}^{c}$ manifold of dimension $n$ and index $r$. The bundle $P_{\text{spin}^{c}}$ is a $\text{spin}^{c}(r,n-r)$-principal bundle, where $\text{spin}^{c}(r,n-r):=\text{spin}(r,n-r)\times_{\mathbb{Z}/2\mathbb{Z}}S^{1}$ and where we write $\text{spin}(r,n-r)$ to denote the connected component of the identity of the spin group sitting inside the Clifford algebra $C\ell_{r,n-r}$. Letting $\delta:\text{spin}(r,n-r)\to\mathrm{GL}(\Sigma)$ be the restriction of an irreducible representation of the complexified Clifford algebra $\mathbb{C}\ell_{r,n-r}$, we define $\Delta:\text{spin}^{c}(r,n-r)\to\mathrm{GL}(\Sigma)$ via $\Delta([s,z])v=z\delta(s)v$. The associated vector bundle $S:=P_{\text{spin}^{c}}\times_{\Delta}\Sigma$ is called the complex spinor bundle $S\to M$, which is a vector bundle of complex dimension $2^{\lfloor\frac{n}{2}\rfloor}$. When $n$ is even, $S$ splits as a direct sum $S=S^{+}\oplus S^{-}$.
@@
where $\{e_{1},\ldots,e_{n}\}$ is a pseudo-orthonormal frame with $\varepsilon_{k}=g(e_{k},e_{k})$ and $i\omega\in\Omega^{2}(M,i\mathbb{R})$ denotes the curvature 2-form coming from the connection 1-form $A\in\Omega^{1}(P_{S^{1}},i\mathbb{R})$ [17, Page 65].
@@ 4.1 Non-local splittings without completeness assumptions
Fix some uniform pair $(I,L_{p}^{\prime})$ for $\Phi$. First, assume that $\Phi|_{I\times L_{p}^{\prime}}$ is injective. Take some point $q\in L_{p}^{\prime}$, with $q\neq p$. Since $L_{p}^{\prime}$ is connected, there exists a path with embedded image $\ell\cong(0,1)$ in $L_{p}^{\prime}$, containing both the point $p$ and the point $q$. Take a tubular neighborhood $\mathcal{U}$ of $\ell$ in $L_{p}^{\prime}$. Since $V$ is nowhere tangent to $\mathcal{U}$ and $\Phi|_{I\times\hskip 0.85358pt\mathcal{U}}$ is injective, we have that $\Phi|_{I\times\hskip 0.85358pt\mathcal{U}}$ is a diffeomorphism onto an open submanifold $N$ of $M$ [32, Theorem 9.20]. Since $N$ has trivial de Rahm cohomology, we can write $V|_{N}=\operatorname{grad}h$, for some function $h:N\to\mathbb{R}$ and note that the intersection of each leaf $L_{m}\in\mathcal{F}$ with $N$ is either empty or a level set of $h$. Hence, if we show that $h(\Phi_{t}(p))=h(\Phi_{t}(q))$, for all $t\in I$, then $\Phi_{t}(q)\in L_{\Phi_{t}(p)}$ and we could conclude the current case. By assumption, it is true for $t=0$. Therefore, it suffices to show that $\frac{d}{dt}h(\Phi_{t}(p))=\frac{d}{dt}h(\Phi_{t}(q))$ for all $t\in I$. We compute that
@@ 5.1 The semi-Riemannian $\text{spin}^{c}$ setting
Case 1 Case 2 Case 3 Case 4
$\frac{i^{r+1}\beta}{2\mu}$ $\frac{i^{r}\beta}{2\mu}$ $\frac{\beta}{2\mu}$ $\frac{\beta}{2\mu}.$
@@
In the Lorentzian case, we have stronger results due to Bohle [11, Section 5] and Leitner [34, Theorem 5.3]. In the Riemannian case, the global statement is essentially the same as due to Baum in [7]. The author is not aware of similar local statements in the Riemannian case but is sure that they are known to experts. There are also stronger results for $r=2$, due to Shafiee and Bahrampour [38]. In case of parallel spinors we have stronger results in [16, 10, 39, 27].
@@ 5.2.1 Type I imaginary generalized Killing spinors
In what follows, we will apply the $\text{spin}^{c}$ Gauß formula from [36, Section 3]. We follow the setup of that paper. Given a leaf $L_{p}\in\mathcal{F}$, we equip it with the induced $\text{spin}^{c}$ structure and we write $S_{L_{p}}$ to denote the induced spinor bundle on the hypersurface $L_{p}$. When the dimension $n$ of $M$ is odd, we have that $S_{L_{p}}\cong S|_{L_{p}}$ and when $n$ is even, we have that $S_{L_{p}}\cong S^{+}|_{L_{p}}$. Hence, when $n$ is odd, we can simply write $\psi|_{L_{p}}$ to denote the restriction of a section, but when $n$ is even we write $\psi|_{L_{p}}$, to mean $\psi^{+}|_{L_{p}}$. In this setup, the $S^{1}$-bundle associated with the $\text{spin}^{c}$ structure on $L_{p}$ is simply the restriction $P_{S^{1}}|_{L_{p}}$, where $P_{S^{1}}$ is the $S^{1}$-bundle associated to the $\text{spin}^{c}$ structure on $M$. Hence, given a connection 1-form $A\in\Omega^{1}(P_{S^{1}},i\mathbb{R})$, we get induced connections $\nabla^{S}$ on $S$ and $\nabla^{S_{L_{p}}}$ on $S_{L_{p}}$. If we let $\bullet:T(L_{p})\otimes S_{L_{p}}\to S_{L_{p}}$ denote the Clifford multiplication on $S_{L_{p}}$, we have that $X\bullet\eta=\tilde{E}\cdot X\cdot\eta$, where on the right hand side, $\eta$ is either viewed as an element of $S|_{L_{p}}$ or $S^{+}|_{L_{p}}$, depending on if $n$ is odd or even.
@@ 5.2.2 Type II imaginary generalized Killing spinors
Recall that an imaginary generalized Killing spinor on a Riemannian $\text{spin}^{c}$ manifold is said to be of type II if the constant $Q_{-}=\langle{\varphi,\varphi}\rangle^{2}-g(V_{\varphi},V_{\varphi})$ is positive. In case the type II imaginary generalized Killing spinor is a Killing spinor, then Große and Nakad [23, Proposition 4.3] showed that the manifold is isometric to hyperbolic space, provided it is complete. It was shown by Rademacher [37] that on spin manifolds, all type II special imaginary generalized Killing spinors have constant Killing function, i.e. they are Killing spinors. Grosse and Nakad [23, Section 4.2.2] extends this to the $\text{spin}^{c}$ case for dimension at least $3$ and provide a counterexample in dimension $2$.